OpenStudy (anonymous):

the longest side of a triangle is 2 times the shortest side. If the third side is 5 more than the shortest side, and the perimeter is 65cm, find the measures of the three sides of the triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

what i needed is the way to set up the equation

OpenStudy (anonymous):

let the smallest side=x then largest side =2x third side= x+5 perimeter P=sum of all sides x+2x+x+5=65 solve for x and then find the sides.